# Diffusion in SPAD Signals

### v1

@@ 3.4 Forward Model
The probability that no photons arrived in the remaining exposure time is derived by Eq. (20)
@@
Eqs. (28, 29, 30) form the PDF, or likelihood, to have a set of detection events described by $(N,\{t_{i}\}_{i=1}^{N})$, given $\lambda$:
@@
To verify that Eq. (31) constitutes a valid PDF, it must satisfy two conditions: non-negativity and normalization.
The first condition, $p(N,\{t_{i}\}|\lambda)\geq 0$, is satisfied by definition. Since the rate parameter $\lambda$ is positive and the exponential function is strictly positive for all real arguments, the likelihood function, constructed as a product of these terms, is strictly positive.
@@ 3.5 Case I
As shown in Fig. 3, after the last measured event $t_{N}$ and its corresponding dead-time period $\tau_{d}$, there is still time left until the exposure time $T$ is exhausted. Therefore,
@@
Using Eq. (15),
@@ 3.6 Case II
Here, after the final measured event $t_{N}$, the time remaining before exhaustion of the exposure time $T$ is shorter than $\tau_{d}$. So,
@@
Integrating over $\{u_{i}\}_{i=1}^{N}$ to calculate $p(N)$ under the integral bounds $0<u_{1}<u_{2}<\dots<u_{N}<[T-(N-1)\tau_{d}]$, with the same methodology as in Eqs. (41, 42),
@@ 3.7 Combined Case
Using Eqs. (54, 55, 56, 57) then proves Eq. (32)

### v2

@@ 3.4 Forward Model
The probability that no detection events happened in the remaining exposure time is derived by Eq. (20)
@@
The probability no detection event happenes at all during exposure time $T$ is also derived by Eq. (20)
$\exp\left[-\lambda T\right]\;.$ (31)
Eqs. (28, 29, 30) form the PDF, or likelihood, to have a set of detection events described by $(N,\{t_{i}\}_{i=1}^{N})$, given $\lambda$ and $N>0$:
@@
In the instance of no detection events $N=0$, the likelihood is described by Eq. (31)
$p(N=0|\lambda)=\exp\left[-\lambda T\right]\;.$ (33)
To verify that Eqs. (32, 33) constitute a valid PDF, they must satisfy two conditions: non-negativity and normalization.
The first condition,
$\begin{split}p(N,\{t_{i}\}_{i=1}^{N}|\lambda)&\geq 0,\quad N\in\{1,2,\dots\}\\ p(N=0|\lambda)&\geq 0\end{split}$ (34)
is satisfied by definition. Since the rate parameter $\lambda$ is positive and the exponential function is strictly positive for all real arguments, the likelihood function, constructed as a product of these terms, is strictly positive.
@@ 3.5 Case I
Assuming at least one detection event $N>0$. As shown in Fig. 3, after the last measured event $t_{N}$ and its corresponding dead-time period $\tau_{d}$, there is still time left until the exposure time $T$ is exhausted. Therefore,
@@
We note that for $N=0$, Eq. (47) coincides with Eq. (33)
$\frac{[\lambda(T-0\cdot\tau_{d})]^{0}\exp[-\lambda(T-0\cdot\tau_{d})]}{0!}=\exp{[-\lambda T]}\;.$ (48)
Therefore, Eq. (47) is true for $N\in\{0,1,2,\dots\}$. Using Eq. (15),
@@ 3.6 Case II
Assuming $N>0$. After the final measured event $t_{N}$, the time remaining before exhaustion of the exposure time $T$ is shorter than $\tau_{d}$. So,
@@
Integrating over $\{u_{i}\}_{i=1}^{N}$ to calculate $p(N)$ under the integral bounds $0<u_{1}<u_{2}<\dots<u_{N}<[T-(N-1)\tau_{d}]$, with the same methodology as in Eqs. (44, 45),
@@
In the instance of no detection events $N=0$, there is no dead time $\tau_{d}$ affecting the model. Therefore, the situation described by $\tt{case\ II}$ cannot occur and
p(N=0)=0.p(N=0)=0\;. (57)
We note that $F_{\text{Erlang}}[T;N=0]$ and $F_{\text{Erlang}}[T+\tau_{d};N=0]$ are the probabilities to detect at least 0 events by times $T$ and $T+\tau_{d}$ respectively:
$\begin{split}F_{\text{Erlang}}&(T;N=0)\\ &=1-\cancelto{0}{\sum_{i=0}^{0-1}}\frac{(\lambda T)^{i}\exp[-\lambda T]}{i!}=1\\[10.0pt] F_{\text{Erlang}}&(T+\tau_{d};N=0)\\ &=1-\cancelto{0}{\sum_{i=0}^{0-1}}\frac{[\lambda(T+\tau_{d})]^{i}\exp[-\lambda(T+\tau_{d})]}{i!}=1\;.\end{split}$ (58)
Assigning $N=0$ in Eq. (56) and using Eq. (58) gives
$\begin{split}p(N=0)&=F_{\text{Erlang}}(T+\tau_{d};N=0)-F_{\text{Erlang}}(T;N=0)\\ &=1-1=0\;,\end{split}$ (59)
which coincides with Eq. (57). Therefore, Eq. (56) is true for $N\in\{0,1,2,\dots\}$.
@@ 3.7 Combined Case
Using Eqs. (61, 62, 63, 64) then proves Eq. (35)